A -10.0 nC point charge and a +20.0 nC point charge are 15.0 cm apart on the x-axis.

A -10.0 nC point and a +20. 0 nC point charge are
Answer #1

Given data:

charges are = -10nC and + 20 nC

distance between them r = 15 cm

a) U=k.Q1.Q2/r. =(9*10^9)(+20*10^-9)(-10*10^-9)/(0.15) = -1.2*10^-5 J

b) F=k*Q1*Q2/r^2
=(9*10^9)(20*10^-9)(10*10^-9)/(0.15)^2
=8*10^-5 N attractive force

c)From negative to the possitive change.